# DiLA: Enhancing LLM Tool Learning with Differential Logic Layer

### v3

@@ 3.4 Differential Logic Layer
Backward Pass. The backward pass is responsible for computing the gradients of the layer inputs and derives updates to variables that steer towards satisfying the constraints $\phi$. A crucial aspect of the backward pass is identifying the input variables that contribute most to the unsatisfiability of the constraint formulas. It is well-established that variables in $\bar{I}$ form the unsatisfiable subset and are, therefore, more likely to be sources of conflict. Conversely, variables not present in $\bar{I}$ can have their gradients set to zero, as their absence in the conflict clauses provides no evidence regarding the correctness or incorrectness of their values. Inspired by the stochastic local search (SLS) algorithm, commonly used in constraint satisfaction problems (Chu et al., 2023), we select the variable with the largest absolute gradient from the candidate set $\bar{I}$ and update its value at each iteration. However, our logic layer diverges from SLS in that it employs a “differential” variable selection mechanism during backpropagation, whereas SLS relies on meta-heuristics.
Algorithm 2 illustrates our backward pass. The backward pass begins by initializing the gradient to zero for all variables (line 1). If $y^{k}$ is false, indicating the presence of unsatisfied clauses, we obtain the set of variables $\bar{I}$ that are present in the falsified clauses $\bar{\phi^{\prime}}$. Once we have obtained the candidate set $\bar{I}$ (line 5), we proceed to select the best variable from this set based on its gradient. Specifically, we compute the gradient as in Equation 5 for each variable in the candidate set (line 7). Then, the logic layer selects a variable and updates its value based on two situations: (1) If there exists a variable with a non-zero gradient (i.e., $g_{i}\neq 0$), the variable with the largest absolute $g_{i}$ would be selected (line 10); (2) If there is no variable satisfying the above condition, indicating that the search is stuck in a local optimum, we randomly select a variable from a falsified clause (line 12) and artificially assigns a gradient that would change its sign after gradient descent (line 13). This criterion guides the updates towards satisfying more clauses at each iteration, as selecting the variable with the largest absolute gradient pushes the loss quantity to decrease in the steepest direction. The forward and backward propagation iterates until the logic layer finds a satisfying assignment or exceeds the maximum time limit.
@@
0:  $\boldsymbol{v}^{k}\in\mathbb{R}^{n}$ from forward pass, $y^{k}$ from forward pass, learning rate $\lambda$.
@@
4 Experiments
Table 1: Accuracy and runtime (s) of CoT, including GPT-4 and Llama-3, LOGIC-LM, and DiLA on LogicalDeduction, simple Boolean Satisfiability and 3-Coloring datasets.
Problem Number of Variables GPT-4 Llama-3 LOGIC-LM DiLA
Acc.(%) Acc.(%) Acc.(%) Time(s) Acc.(%) Time(s)
Logical
Deduction
3,5,7 76 81 100 0.01 100 0.01
SAT 20 12 17 100 0.05 100 0.01
50 5 0 100 0.12 100 0.01
100 0 0 100 0.17 100 0.03
200 0 0 100 0.99 100 0.04
250 0 0 100 5.24 100 0.08
Graph Coloring 10 20 22 100 0.05 100 0.01
50 7 10 100 0.19 100 0.02
100 0 0 100 0.58 100 0.04
150 0 0 100 2.64 100 0.15
200 0 0 100 5.70 100 0.24
4.1 Setup
Tasks: We conduct experiments on three fundamental reasoning tasks: the logical deduction problem, Boolean Satisfiability, and Graph Coloring problems. The logic deduction problems are mostly about deducing the order of a sequence of objects from a minimal set of conditions. Here we utilize LogicalDeduction dataset from the BigBench (Srivastava et al., 2022) collaborative benchmark. For Boolean Satisfiability problems, we utilize open-source benchmark instances  [^2] https://www.cs.ubc.ca/~hoos/SATLIB/benchm.html with 20 to 250 variables, focusing on finding variable assignments that satisfy all constraints. Notably, we do not consider the unsatisfiable cases as finding unsatisfied certificates is a distinct problem. For Graph Coloring problems, we randomly generated 100 3-coloring instances with vertices counts ranging from 10 to 200, following the approach in  (Stechly et al., 2024) (see Appendix C for details), aiming to color graph vertices so that no two adjacent vertices share the same color. Furthermore, we also incorporate several complex SAT cases in SAT Competition 2023333https://satcompetition.github.io/2023/ and open-source graph coloring problems with large number of edges [^4] https://www.cs.ubc.ca/~hoos/SATLIB/Benchmarks/SAT/ to illustrate the robustness of our proposed DiLA. These instances originate from practical industrial problems and frequently pose challenges for modern SAT solvers. Notably, current solvers often enter into an infinite loop when handling these test cases due to their complex structure, requiring a prohibitively long time to solve.
Baselines: We conducted a comparative analysis between DiLA and three baselines, including the prompt-based methods with GPT-4 (Achiam et al., 2023) and Llama-3 (Llama-3-70B-Instruct) (Meta, 2024), and the solver-augmented method, LOGIC-LM (Pan et al., 2023). In particular, LOGIC-LM (Pan et al., 2023) employs GPT-4 to parse problem specifications and offloads the logical reasoning task to the symbolic solvers, including CSP Solver [^5] https://github.com/python-constraint/ python-constraint and Z3 solver (De Moura and Bjørner, 2008), which serves as the state-of-the-art logic reasoning methods.
Setup: We implement a prototype of our proposed DiLA using Pytorch (Paszke et al., 2019), leveraging GPT-4 as the backbone LLM model. Notably, the logic layer within DiLA has no training parameters and can adapt to various problem types expressed in Boolean variables. We employ the Adam optimizer (Kingma and Ba, 2014) with a learning rate of $2\times 10^{-1}$ in DiLA, facilitating the effective updating of selected variables. Furthermore, we use a temperature of 0 for LLMs, consistent with the LOGIC-LM approach. We set a time limit of 10,000 seconds for both the solvers and the logic layer.
4.2 Main Results
We report the accuracy of DiLA and the three baselines in Table 1. Accuracy is evaluated based on whether the LLM can output a correct answer that satisfies all constraints. We evaluate LLMs over 100 instances in each domain. In addition to accuracy, we also report the solving runtime for LOGIC-LM and DiLA, both of which leverage GPT-4 as a backbone LLM to perform language understanding. Analysis of Table 1 reveals that both solver-augmented LOGIC-LM and logic layer-aided DiLA achieve 100% accuracy on these simple benchmarks, owing to the precise logic parsing and accurate solving. In contrast, standalone LLMs, including GPT-4 and Llama-3, cannot handle reasoning problems with over 100 variables with CoT prompting. More importantly, for all test instances, DiLA exhibits faster performance than LOGIC-LM in solving process, especially for relatively large cases with over 200 variables. The runtime speedup can be up to 65.5$\times$ when dealing with SAT problems with 250 variables and 23.8$\times$ when tackling 3-coloring problems with 200 nodes. Overall, both the solver-augmented LLM and our proposed DiLA can successfully address these simple artificial test cases, with the runtime speedup highlighting the efficiency of DiLA in logic reasoning.
Table 2: Runtime (s) of Z3, Kissat, and our DiLA on solving hard reasoning problems.
@@ 4.2 Main Results
4.3 Robustness of Reasoning
Compared with SATLM (Ye X, 2023) and LOGIC-LM (Pan et al., 2023), which rely on an off-the-shelf solver for reasoning, DiLA sidesteps the limitations of the solvers themselves, such as high computational costs when dealing with complex formulas. In practice, there are intricate cases where even state-of-the-art SAT solvers struggle, sometimes taking weeks to solve. These challenging instances often involve Boolean satisfiability formulas with complex structures and graph coloring problems with considerable edge connections. For our evaluation, we leverage the state-of-the-art SMT solver, Z3 (De Moura and Bjørner, 2008), and the widely-used SAT solver, Kissat (Biere and Fleury, 2022), as baselines, and test on some hard test cases derived from real-world problems.
@@ 4.3 Robustness of Reasoning
The experimental results are shown in Table 2, revealing several key observations. First, DiLA successfully solves these challenging cases within a reasonable runtime, whereas both Z3 and Kissat fail to produce valid results within the time limit. As indicated in Table 2, these difficult constraint satisfaction instances typically exhibit large clause-to-variable (CV) ratios (m/n). For example, the test case “rbsat” consists of only 1150 variables, but the total number of clauses amounts to 84,314, resulting in an extremely high CV ratio of 73.32. This suggests that there is likely only one viable solution for these problems. If a traditional solver’s initial search path deviates significantly from the correct path, the CDCL framework’s inability to rectify errors through a learning-from-mistakes system can lead to an endless loop. In contrast, our proposed DiLA, initialized by LLMs and guided by differentiation of the loss function, enables us to first reach a partially satisfied solution and then progressively update it using an efficient searching strategy, rather than completely failing if stuck.
@@
Second, the initial solution generated by backbone LLMs can serve as an excellent starting point. For example, in the SAT case “SCPC”, after analyzing all constraints, the backbone LLM provides an initial solution with an unusual all-false variable assignment. We discovered that over 99% of the final satisfying variable assignments should indeed be set to 0, indicating that a significant proportion of variables require no further updates. Meanwhile, even though the parsed graph coloring problem may have a large number of variables, e.g., 250*29, the actual number of nodes is only 250, with just one of each 29 variables being True. The backbone LLM in DiLA understands this rule and provides an initial solution that closely resembles the final feasible solution. In contrast, traditional solvers rely on random initialization and need to explore all possible assignments for each variable before reaching the final solution, resulting in exponential search spaces and reduced efficiency. In general, the evaluation results on real-world cases demonstrate the high efficiency of our proposed DiLA.
5 Related Works
Prior approaches to NL-based reasoning with LLMs can be broadly categorized into two groups. One is in-context learning approaches that design special prompts to elicit LLMs’ step-by-step reasoning capabilities. Typical methods include chain-of-thought prompting (Wang et al., 2023) that generates a sequence of reasoning steps before the final answer and the least-to-most prompting (Zhou et al., 2022) that breaks the problem down into simpler components that can be solved sequentially. Both the above approaches perform reasoning directly over natural language (NL), providing greater flexibility than symbolic-based reasoning. However, the intrinsic complexity and ambiguity of NL also bring undesired issues such as unfaithful reasoning and hallucinations. The other is tool-augmented approaches that only require LLMs to parse the problem specification out of the language description accurately and leverage off-the-shelf automated tools to derive the final answer, as exemplified by SATLM (Ye X, 2023) and LOGIC-LM (Pan et al., 2023). The tool-augmented approaches guarantee the correctness of the answer with respect to the parsed specification and avoid planning errors in the solving process. However, the performance of such tool-augmented methods highly relies on external tools and can not deal with many real problems due to the deficiency of backbone solvers.
6 Conclusion
@@
In this work, we introduce a pioneering method named differential logic layer-aided language modeling (DiLA). Starting with an NL reasoning problem, DiLA first uses an LLM to cast it into a SAT problem and generate a possible solution based on its language understanding, and then progressively refines this solution within a logic layer. In this way, we harness the potential of the language-understanding ability of LLMs and sidestep the limitations of off-the-shelf solvers. Extensive experiments on two reasoning tasks demonstrate the superior efficiency of our approach over state-of-the-art solver-augmented LLMs. Notably, DiLA demonstrates its robustness in dealing with challenging problem instances that SOTA solvers struggle with. The results suggest that DiLA achieves new state-of-the-art performance in symbolic logical reasoning tasks, paving the way for more applications of LLMs in practical reasoning settings.
7 Limitation
We identify two main limitations of DiLA. First, DiLA relies on translating reasoning problems into logical formats that can be encoded by boolean variables and first-order logic. As a consequence, the model’s applicability is inherently bounded by the expressiveness of the underlying theory, as not all problems can be easily encoded in first-order formulas with boolean vectors. Nevertheless, the wide range of tasks that we can instantiate our DiLA framework on shows its general applicability.
Second, DiLA relies on backbone LLMs to translate natural language problems into symbolic representations. Although this approach has shown promise, it may struggle with logical representations featuring complex grammar structures, such as probabilistic circuit logic. This limitation stems from the challenge of communicating intricate grammatical rules to the language model through a limited number of demonstrations within a constrained context size. To address this, future research could investigate the development of specialized modules to improve the mapping between natural language and symbolic language, for instance, by fine-tuning LLMs with symbolic formulas for specific reasoning tasks.
Appendix A Details of the SAT Specification

### v4

@@ Introduction
Table 1: Accuracy and runtime (s) of CoT, including GPT-4, Llama-3, and DeepSeek-R1, SATLM, and DiLA on LogicalDeduction, Boolean Satisfiability and 3-Coloring datasets. Llama-3 SC denotes the Llama-3 model with Self-Consistency (Wang et al. 2022).
Problem #Variables GPT-4 Llama-3 Llama-3 SC DeepSeek-R1 SATLM DiLA
Acc (%) Acc (%) Acc (%) Acc (%) Acc (%) Time (s) Acc (%) Time (s)
Logical
Deduction
3,5,7 76 81 83 97 100 0.01 100 0.01
SAT 20 12 17 17 39 100 0.05 100 0.01
50 5 0 0 7 100 0.12 100 0.01
100 0 0 0 0 100 0.17 100 0.03
200 0 0 0 0 100 0.99 100 0.04
250 0 0 0 0 100 5.24 100 0.08
Graph
Coloring
10 20 22 22 34 100 0.05 100 0.01
50 7 10 10 15 100 0.19 100 0.02
100 0 0 0 0 100 0.58 100 0.04
150 0 0 0 0 100 2.64 100 0.15
200 0 0 0 0 100 5.70 100 0.24
Backward Pass. The backward pass is responsible for computing the gradients of the layer inputs and derives updates to variables that steer towards satisfying the constraints $\phi$. A crucial aspect of the backward pass is identifying the input variables that contribute most to the unsatisfiability of the constraint formulas. It is well-established that variables in $\bar{I}$ form the unsatisfiable subset and are, therefore, more likely to be sources of conflict. Conversely, variables not present in $\bar{I}$ can have their gradients set to zero, as their absence in the conflict clauses provides no evidence regarding the correctness or incorrectness of their values. Inspired by the stochastic local search (SLS) algorithm, commonly used in constraint satisfaction problems (Chu, Cai, and Luo 2023), we select the variable with the largest absolute gradient from the candidate set $\bar{I}$ and update its value at each iteration. However, our logic layer diverges from SLS in that it employs a “differential” variable selection mechanism during backpropagation, whereas SLS relies on meta-heuristics.
@@
[…]
@@
Algorithm 2 illustrates our backward pass. The backward pass begins by initializing the gradient to zero for all variables (line 1). If $y^{k}$ is false, indicating the presence of unsatisfied clauses, we obtain the set of variables $\bar{I}$ that are present in the falsified clauses $\bar{\phi^{\prime}}$. Once we have obtained the candidate set $\bar{I}$ (line 5), we proceed to select the best variable from this set based on its gradient. Specifically, we compute the gradient as in Equation 5 for each variable in the candidate set (line 7). Then, the logic layer selects a variable and updates its value based on two situations: (1) If there exists a variable with a non-zero gradient (i.e., $g_{i}\neq 0$), the variable with the largest absolute $g_{i}$ would be selected (line 10); (2) If there is no variable satisfying the above condition, indicating that the search is stuck in a local optimum, we randomly select a variable from a falsified clause (line 12) and artificially assign a gradient that would change its sign after gradient descent (line 13). More discussions are given in Appendix.
Experiments
In this section, we present an empirical evaluation of DiLA on solving logical reasoning problems expressed by natural language. Particularly, we test our approach in satisfiable datasets, as tackling unsatisfiable (UNSAT) certification presents a distinct and separate challenge. Although DiLA is not directly applicable to UNSAT problems, we present an extension that identifies potential unsatisfiable cores in AppendixB. Many real-world applications (e.g., resource allocation, route planning) naturally admit solutions when given sufficient flexibility. For instance, in graph coloring, while a graph may be UNSAT for 3 colors, it often becomes satisfiable when more colors are permitted. For satisfiable instances, LLMs need to determine the CNF instance’s satisfiability and output the variable assignments that genuinely satisfy the input CNF instances.
Table 2: Runtime (s) of Z3, Kissat, and DiLA on hard problems. DiLA successfully handles these hard cases where canonical solvers struggle.
@@
sgen3g250.29SCPC0300300150150107.33107.33269.65269.65137.22137.2211.8311.8330.0730.0720.2320.23Runtime(s)\!~(s)w/o. LLMw. LLM
Figure 4: Comparing DiLA with and without LLM initialization.
Experiment Setup
Tasks. We conduct experiments on three fundamental reasoning tasks: the logical deduction problem, Boolean Satisfiability, and Graph Coloring problems. The logic deduction problems are mostly about deducing the order of a sequence of objects from a minimal set of conditions. Here we utilize the LogicalDeduction dataset from the BigBench (Srivastava et al. 2022) collaborative benchmark. For SAT problems, we utilize open-source benchmark instances  [^2] https://www.cs.ubc.ca/~hoos/SATLIB/benchm.html with 20 to 250 variables, focusing on finding variable assignments that satisfy all constraints. For Graph Coloring problems, we randomly generated 100 3-coloring instances with vertex counts ranging from 10 to 200, following the approach in  (Stechly, Valmeekam, and Kambhampati 2024) (see AppendixE for details), aiming to color graph vertices so that no two adjacent vertices share the same color. Furthermore, we also incorporate several complex SAT cases in SAT Competition 2023 and open-source graph coloring problems with a large number of edges to illustrate the robustness of our proposed DiLA. These instances that originate from practical industrial problems are known to be satisfiable, reflecting real-world scenarios where solutions are expected to exist. They frequently pose challenges for modern SAT solvers. Notably, current solvers often enter into an infinite loop when handling these test cases due to their complex structure, requiring a prohibitively long time to solve.
Baselines. We conducted a comparative analysis between DiLA and several baselines, including the CoT-based methods with GPT-4 (Achiam et al. 2023) and Llama-3 (Llama-3-70B-Instruct) (Meta 2024), Self-Consistency (Wang et al. 2022) enhanced Llama-3, DeepSeek-R1 (Guo et al. 2025), and the solver-augmented method, SATLM (Ye X 2023). In particular, SATLM (Ye X 2023) employs GPT-4 to parse problem specifications and offloads the logical reasoning task to the symbolic solvers, i.e., Z3 solver (De Moura and Bjørner 2008), which serves as the state-of-the-art tool-learning method for logic reasoning.
Setup. We implement a prototype of our proposed DiLA using Pytorch (Paszke et al. 2019), leveraging GPT-4 as the backbone LLM model. Notably, the logic layer within DiLA has no training parameters and can adapt to various problem types expressed in Boolean variables. We employ the SGD optimizer with a learning rate of $2\times 10^{-1}$ in DiLA, facilitating the effective updating of selected variables. Furthermore, we use a temperature of 0 for LLMs, consistent with the SATLM approach. We set a time limit of 10,000 seconds for both the solvers and the logic layer. All experiments are performed on 8 NVIDIA V100 (32GB), with LLM-involved stages dominating memory usage. The logic layer of DiLA can be performed on a single GPU for benchmarks with hundreds of thousands of variables/clauses, due to sparse connections and low-bit layer weights.
Main Results
We report the accuracy of DiLA and baselines in Table 1. Accuracy is evaluated based on whether the LLM can output a correct answer that satisfies all constraints. We evaluate LLMs over 100 instances in each domain. In addition to accuracy, we also report the solving runtime for SATLM and DiLA, both of which leverage GPT-4 as a backbone LLM to perform language understanding.
[…]
Generalization to Industrial Benchmarking
To investigate the performance boundaries and demonstrate the differentiability of DiLA from solver-aided methods, we evaluate it on a set of hard problems derived from the real world. We compare DiLA against two distinct and powerful solvers: Z3 and Kissat. For all methods, we set a stringent time limit of 10,000 seconds per instance. All experiments were conducted on the same hardware configuration as previously described.
@@
The experimental results, presented in Table 2, demonstrate that DiLA successfully solves these challenging cases within a reasonable runtime, whereas both Z3 and Kissat fail to produce valid results within the time limit. As indicated in Table 2, these difficult constraint satisfaction instances typically exhibit large clause-to-variable (CV) ratios (m/n). For example, the test case “rbsat” consists of only 1150 variables, but the total number of clauses amounts to 84,314, resulting in an extremely high CV ratio of 73.32. This suggests that there is likely only one viable solution for these problems. If a traditional solver’s initial search path deviates significantly from the correct path, the CDCL framework’s inability to rectify errors through a learning-from-mistakes system can lead to an endless loop. In contrast, our proposed DiLA, initialized by LLMs and guided by differentiation of the loss function, enables us to first reach a partially satisfied solution and then progressively update it using an efficient searching strategy, rather than completely failing if stuck. In general, the evaluation results on industrial cases demonstrate the high efficiency of our proposed DiLA.
@@
Ablation study of LLM initialization. The comparison results, shown in Figure 4, indicate that the initial solutions generated by backbone LLMs can serve as an excellent starting point. Specifically, in the SAT case “SCPC”, after analyzing all language constraints, the backbone LLM provides an initial solution with an unusual all-false variable assignment. We discovered that over 99% of the final satisfying variable assignments should indeed be set to 0, indicating that a significant proportion of variables require no further updates. In contrast, traditional solvers rely on random initialization and need to explore all possible assignments for each variable before reaching the final solution, resulting in exponential search spaces and reduced efficiency.
Natural Language Reasoning
While previous sections demonstrated DiLA’s effectiveness on standard benchmarks, we now evaluate its performance on a more challenging task: solving constraint satisfaction problems described entirely in natural language, without any structured inputs. This evaluation directly addresses the concern about the parsing capability of the backbone LLM and demonstrates DiLA’s end-to-end reasoning capability from raw text to validated solutions.
We construct a new benchmark, Natural Language Constraint Reasoning (NLCR), specifically designed to test the synergy between language understanding and logical reasoning. Each problem in NLCR requires three key capabilities: (1) extracting implicit constraints from unstructured natural language text, (2) translating semantic constraints into formal SAT specifications, and (3) finding satisfying variable assignments that meet all requirements.
The benchmark comprises 100 problems across four realistic domains: scheduling problems (30 instances) involving temporal and resource constraints, resource allocation (25 instances) with capacity and assignment rules, planning problems (25 instances) requiring sequential decision-making, and configuration problems (20 instances) with dependency constraints. An instance from NLCR benchmark is illustrated in AppendixA.
Comparative System Evaluation
We compare three representative approaches that span from pure neural to hybrid neuro-symbolic methods: (1) Pure LLM leveraging GPT-4 with chain-of-thought prompting to directly generate solutions, (2) SATLM combining LLM parsing with the Z3 symbolic solver for formal reasoning, and (3) DiLA integrating LLM parsing and initialization with our differential logic layer for iterative refinement.
Table 3 presents the overall performance across all 100 NLCR problems. DiLA achieves 87% end-to-end success rate, outperforming SATLM (76%) and Pure LLM (51%) by significant margins. Notably, DiLA also demonstrates superior efficiency with an average solving time of 4.5 seconds, faster than SATLM’s 8.7 seconds despite handling more complex reasoning internally. The parsing accuracy of DiLA (92%) exceeds both baselines, suggesting that the iterative refinement capability of the logic layer provides implicit feedback that improves constraint extraction quality.
Table 3: Performance comparison on NLCR benchmark
System Parse Solution E2E Time
Acc Quality Succ. (s)
Pure LLM 78% 65% 51% 3.2
SATLM 85% 90% 76% 8.7
DiLA 92% 95% 87% 4.5
Breaking down performance by problem category in Table 4, we observe that DiLA’s advantages are most pronounced in scheduling tasks (+17% over SATLM), where temporal reasoning and resource constraints interact in complex ways. The improvement is consistent across resource allocation (+12%) and planning problems (+12%), while configuration problems show comparable performance. This pattern suggests that DiLA’s strength lies in handling problems with intricate constraint interactions that challenge both pure LLM reasoning and traditional solver heuristics.
Table 4: End-to-end success rate by problem category.
Category LLM SATLM DiLA Gain
Scheduling 47% 73% 90% +17%
Resource 52% 76% 88% +12%
Planning 48% 72% 84% +12%
Config. 60% 85% 85% 0%
Overall 51% 76% 87% +11%
Quantifying Semantic Understanding
Beyond overall performance metrics, we investigate whether DiLA’s LLM component provides value beyond constraint parsing through semantic understanding. We design 20 problems with explicit semantic biases where understanding the problem structure should guide initialization. For instance, in a star-shaped graph coloring problem with one central node connected to 10 peripheral nodes, the semantic insight is that the central node should receive a “rare” color since it must differ from all peripherals.
We define a Semantic Understanding Score (SUS) as the fraction of variables that receive semantically meaningful initial assignments:
$\text{SUS}=\frac{\text{\# semantically correct initial assignments}}{\text{\# total variables}}$ (6)
Table 5 demonstrates that LLM initialization achieves a SUS of 0.89, correctly handling 18 out of 20 semantic-rich problems. This substantially exceeds random initialization (SUS = 0.18, 2/20 correct) and heuristic-based initialization (SUS = 0.45, 9/20 correct). The high semantic understanding translates directly into faster convergence, as the logic layer starts from a more informed initial state that already captures key problem structure.
Table 5: Semantic Understanding Score across 20 problems.
Method SUS Score Correct
Random 0.18 2/20
Heuristic 0.45 9/20
LLM (DiLA) 0.89 18/20
Conclusion
@@
In this work, we introduce a pioneering method named differential logic layer-aided language modeling (DiLA). Starting with an NL reasoning problem, DiLA first uses an LLM to cast it into a SAT problem and generate a possible solution based on its language understanding, and then progressively refines this solution within a logic layer. In this way, we harness the potential of the language-understanding ability of LLMs and sidestep the limitations of off-the-shelf solvers. Extensive experiments on two reasoning tasks demonstrate the superior efficiency of our approach over state-of-the-art solver-augmented LLMs. On large-scale industrial verification and cryptographic problems, DiLA retains its efficiency advantage where state-of-the-art solvers (Z3, Kissat) fail to converge within their time budgets, underscoring its robustness and scalability. Beyond traditional SAT formulations, DiLA exhibits strong synergy between language understanding and logical reasoning in natural language constraint reasoning tasks, achieving 87% end-to-end success rate on our newly proposed NLCR benchmark. The results suggest that DiLA achieves new state-of-the-art performance in symbolic logical reasoning tasks, paving the way for more applications of LLMs in practical reasoning settings.
Appendix A More Evaluations on Natural Language Reasoning
To test the limits of each approach, we evaluate performance on 10 large-scale NLCR instances with 20–30 entities, 50–80 constraints, and 500–1000 SAT variables after encoding. These problems represent realistic complexity levels encountered in industrial applications.
Figure 5 presents a representative scheduling problem that illustrates the complexity of natural language constraint specification in our benchmark.
Problem ID: NLCR-Schedule-015 Description: A company has 5 meeting rooms (R1–R5) and needs to schedule 8 meetings (M1–M8) during a single day. Each meeting lasts 1 hour, and the working day spans 9 AM to 6 PM (9 time slots). Constraints: 1. Meeting M1 must occur before M3. 2. M2 and M4 cannot overlap. 3. M5 requires room R1 (morning only). 4. M6 and M7 must be adjacent. 5. R3 cannot be used >>2 consecutive hours. 6. M8 must be the last meeting. 7. No room hosts multiple meetings simultaneously. Task: Provide a feasible schedule assigning each meeting to a room and time slot.
Figure 5: Example problem from the NLCR benchmark requiring natural language understanding and constraint reasoning.
Table 6 reveals stark performance differences across systems. Pure LLM fails on all instances (0% success rate), unable to maintain reasoning coherence at this scale due to the depth and breadth of constraint interactions. SATLM achieves 40% success rate, which can be decomposed into its constituent success factors: 60% parsing accuracy multiplied by 67% solver success rate (4 out of 6 correctly parsed instances solved within timeout). The solver timeouts occur primarily on instances with high clause-to-variable ratios where Z3’s CDCL-based search struggles.
[…]
Table 6: Performance on large-scale NLCR problems with 20–30 entities and 50–80 constraints (timeout = 300s).
Problem ID #Entities #Constraints #Variables Pure LLM SATLM DiLA Speedup
[…]
Failure Mode Analysis
To understand the limitations of each approach, we categorize all failures across the 100 NLCR problems in Table 7. The analysis reveals distinct bottlenecks for each system. Pure LLM exhibits a high rate of logical inconsistencies (27%), where the generated solutions violate constraints despite the model’s apparent understanding of individual requirements. This occurs because LLMs lack systematic constraint verification during step-by-step reasoning, leading to solutions that satisfy some constraints while inadvertently violating others.
SATLM’s primary failure mode is solver timeout (17%), occurring when the Z3 solver encounters complex formulas that exceed its heuristic search capabilities within the time limit. An additional 15% of failures stem from parsing errors where the LLM misses implicit constraints or incorrectly formalizes natural language specifications. These parsing failures are particularly problematic for SATLM because the solver has no mechanism to recover from incomplete or incorrect constraint formulations.
DiLA achieves the lowest total failure rate (13%) by mitigating both classes of failures. The logic layer’s gradient-based refinement handles solver-hard instances that cause SATLM timeouts, reducing timeout failures to just 3%. Meanwhile, the iterative nature of DiLA’s reasoning process appears to improve parsing robustness, lowering parsing errors to 8%. The virtual absence of partial solutions (0%) demonstrates that when DiLA succeeds in parsing, the logic layer reliably finds complete satisfying assignments.
Table 7: Failure mode distribution across 100 NLCR problems.
Failure Type LLM SATLM DiLA
Parsing error 22% 15% 8%
Logical inconsistency 27% 3% 2%
Solver timeout 17% 3%
Partial solution 8% 0%
Total 49% 24% 13%
The LLM component in DiLA provides value that extends beyond merely extracting explicit constraints from text. Figure 6 illustrates this through a medical appointment scheduling problem involving both hard constraints (“Alice’s appointment must be after Bob’s”) and soft preferences (“Carol prefers mornings but can do afternoons if necessary”).
Problem: Schedule 3 medical appointments. Alice’s appointment must be after Bob’s. Carol prefers mornings but can do afternoons if necessary. SATLM Parsing: Constraint 1: $t_{\text{Alice}}>t_{\text{Bob}}$ Constraint 2: cannot understand “preference” DiLA with LLM Understanding: Constraint 1: $t_{\text{Alice}}>t_{\text{Bob}}$ (hard) ✓ Constraint 2: Soft – prefer $t_{\text{Carol}}\in\{9,10,11\}$ AM Initial solution: Bob: 9 AM, Alice: 11 AM, Carol: 10 AM (honors preference) Logic layer refines based on hard constraints while preserving soft preferences when possible.
Figure 6: Example showing DiLA’s nuanced understanding of constraint priorities (hard vs. soft) beyond pure logical parsing.
SATLM’s parsing treats “preference” as non-binding and ignores it entirely, leading to solutions that may unnecessarily violate user preferences. In contrast, DiLA’s LLM component recognizes the distinction between hard constraints and soft preferences, encoding this in the initial solution by assigning Carol to a morning slot (10 AM) while ensuring the hard constraint (Alice after Bob) is satisfied. The logic layer then refines this initialization while attempting to preserve the soft preference when compatible with hard constraints.
This example demonstrates three key capabilities of DiLA’s LLM component: (1) nuanced understanding of constraint priorities (hard versus soft), (2) contextual reasoning that goes beyond pure logical parsing to capture user intent, and (3) generation of human-aligned solutions that respect implicit preferences embedded in natural language descriptions.
Details of the SAT Specification